What is the water solubility of BaCO2 (Kp 2.6 x 10 in a solution containing 0.60 M CO (aq)?
Answer #1

Tom Bolog - Ba 2+ + 0. Co, 2- 0.60 Ksp = [Ba2+) {log 2) 2.GX16-9 () (0.60+5) S = 413388-9 m (892+] = 4.33x109 m